OpenStudy (anonymous):

Find the solution: 9x-3y=8

OpenStudy (anonymous):

of x or y?

OpenStudy (anonymous):

I want to know if the solution is (4,-8)

OpenStudy (anonymous):

oh then just plug in 4 for x and -8 for y so you would have 9(4)-3(-8) =8 36+24=8 60=8 it doenst equal, so that means that (4,-8) is NOT a solution

OpenStudy (angela210793):

if u want to know that all u have 2 do is put 4 instead of x and -8 instead of y..if it is=0 then tht's the solution :)

OpenStudy (anonymous):

if x =0 or y =0 then x=8/9, y=0 or x=0, y=8/-3